# RRWNet: Recursive Refinement Network for Effective Retinal Artery/Vein Segmentation and Classification

### v2

@@ 6.3.2 A/V segmentation and classification for all GT vessels and RR post-processing
RITE Morano et al., (2021) 89.26 (94.37) 81.49 (86.46) 97.37 (97.95) 13.71 (28.79) 86.02 (70.84) 87.26 (86.97) 98.12 (98.16) 27.54 (40.95) 72.23 (58.74)
Galdran et al., (2022) 90.59 (94.80) 83.26 (87.21) 97.31 (98.02) 9.93 (27.15) 89.85 (72.48) 87.71 (84.81) 98.21 (98.27) 16.84 (39.14) 82.84 (60.47)
Karlsson and Hardarson, (2022) 94.67 (94.70) 86.39 (86.81) 97.79 (97.62) 14.42 (33.19) 85.30 (66.42) 89.47 (84.47) 98.27 (97.76) 22.39 (42.91) 77.35 (56.75)
Chen et al., (2022) 90.91 (93.93) 80.94 (84.03) 94.81 (96.49) 19.04 (29.84) 80.56 (69.80) 85.75 (85.50) 95.18 (97.64) 25.16 (46.78) 74.67 (52.68)
@@
HRF Morano et al., (2021) 94.76 (96.32) 84.02 (84.64) 98.86 (99.01) 44.87 (51.80) 54.73 (47.87) 87.85 (83.83) 98.96 (99.07) 46.27 (46.13) 53.20 (53.07)
Galdran et al., (2022) 93.94 (96.65) 82.65 (85.26) 98.91 (99.06) 17.07 (55.80) 82.33 (43.40) 86.94 (88.64) 98.76 (99.04) 18.07 (52.07) 81.47 (47.07)
Karlsson and Hardarson, (2022) 95.80 (96.91) 83.27 (82.60) 98.55 (98.52) 31.80 (60.87) 67.93 (38.40) 86.42 (83.68) 98.41 (98.50) 23.60 (54.20) 75.67 (45.00)
Chen et al., (2022) 92.08 (96.72) 77.95 (81.60) 93.75 (98.06) 27.20 (48.87) 72.67 (50.73) 82.12 (82.46) 94.58 (97.99) 36.33 (44.47) 63.53 (54.93)
@@
GT
RRWNet (ours)
Morano et al., (2021)
Chen et al., (2022)
Galdran et al., (2022)
Karlsson and Hardarson, (2022)
@@
GT
RRWNet (ours)
Morano et al., (2021)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Chen et al., (2022)
Galdran et al., (2022)
Karlsson and Hardarson, (2022)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Overall, the segmentation maps generated by RRWNet are more accurate and topologically consistent than the segmentation maps obtained by the other methods. For example, while the other methods tend to mix the classification of a vessel that is difficult to classify, the proposed RRWNet is able to correctly classify the whole vessel as either an artery or a vein (see Fig. 8, blue box). In addition, RRWNet correctly classifies pixels at vessel crossings as belonging to both the artery and vein classes simultaneously (in the figures, represented by white pixels), while the other methods, except Chen et al., (2022), tend to classify them as only one of the two classes or leave them unclassified (with low probability for both classes), leading to discontinuities in the segmentation maps. This RRWNet behavior inherently leads to more topologically consistent segmentation maps.
Additionally, Fig. 10 shows examples of the segmentation maps obtained by the model proposed by Galdran et al., (2022) before and after applying the proposed RR module as a post-processing step.
GT
Galdran et al., (2022)
Galdran et al., (2022) with RR
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Examples of segmentation maps obtained by the model of Galdran et al., (2022) before and after applying the proposed RR module as a post-processing step. A few notable differences between the resulting segmentation maps are highlighted in colored boxes. [RITE, image 11]
As evidenced by the figure, the use of this module demonstrably enhances the quality of the segmentation results. Notably, the RR module effectively addresses the issue of false positive veins in the base segmentation map by accurately reclassifying them as arteries. Additionally, it successfully bridges the gaps between arteries and veins at crossing points (represented in white), where the original segmentation predicted a low probability for one or both classes (represented in dark purple). As mentioned above, this issue is consistently observed in the outputs of other methods, with the exception of Chen et al., (2022).

### v3

@@ 6.3.2 A/V segmentation and classification for all GT vessels and RR post-processing
RITE Morano et al., [2021] 89.26 (94.37) 81.49 (86.46) 97.37 (97.95) 13.71 (28.79) 86.02 (70.84) 87.26 (86.97) 98.12 (98.16) 27.54 (40.95) 72.23 (58.74)
Galdran et al., [2022] 90.59 (94.80) 83.26 (87.21) 97.31 (98.02) 9.93 (27.15) 89.85 (72.48) 87.71 (84.81) 98.21 (98.27) 16.84 (39.14) 82.84 (60.47)
Karlsson and Hardarson, [2022] 94.67 (94.70) 86.39 (86.81) 97.79 (97.62) 14.42 (33.19) 85.30 (66.42) 89.47 (84.47) 98.27 (97.76) 22.39 (42.91) 77.35 (56.75)
Chen et al., [2022] 90.91 (93.93) 80.94 (84.03) 94.81 (96.49) 19.04 (29.84) 80.56 (69.80) 85.75 (85.50) 95.18 (97.64) 25.16 (46.78) 74.67 (52.68)
@@
HRF Morano et al., [2021] 94.76 (96.32) 84.02 (84.64) 98.86 (99.01) 44.87 (51.80) 54.73 (47.87) 87.85 (83.83) 98.96 (99.07) 46.27 (46.13) 53.20 (53.07)
Galdran et al., [2022] 93.94 (96.65) 82.65 (85.26) 98.91 (99.06) 17.07 (55.80) 82.33 (43.40) 86.94 (88.64) 98.76 (99.04) 18.07 (52.07) 81.47 (47.07)
Karlsson and Hardarson, [2022] 95.80 (96.91) 83.27 (82.60) 98.55 (98.52) 31.80 (60.87) 67.93 (38.40) 86.42 (83.68) 98.41 (98.50) 23.60 (54.20) 75.67 (45.00)
Chen et al., [2022] 92.08 (96.72) 77.95 (81.60) 93.75 (98.06) 27.20 (48.87) 72.67 (50.73) 82.12 (82.46) 94.58 (97.99) 36.33 (44.47) 63.53 (54.93)
@@
GT RRWNet (ours) Morano et al., [2021]
Refer to caption Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Chen et al., [2022] Galdran et al., [2022] Karlsson and Hardarson, [2022]
Refer to caption Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
GT RRWNet (ours) Morano et al., [2021]
Refer to caption Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Chen et al., [2022] Galdran et al., [2022] Karlsson and Hardarson, [2022]
Refer to caption Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Overall, the segmentation maps generated by RRWNet are more accurate and topologically consistent than the segmentation maps obtained by the other methods. For example, while the other methods tend to mix the classification of a vessel that is difficult to classify, the proposed RRWNet is able to correctly classify the whole vessel as either an artery or a vein (see Fig. 8, blue box). In addition, RRWNet correctly classifies pixels at vessel crossings as belonging to both the artery and vein classes simultaneously (in the figures, represented by white pixels), while the other methods, except Chen et al., [2022], tend to classify them as only one of the two classes or leave them unclassified (with low probability for both classes), leading to discontinuities in the segmentation maps. This RRWNet behavior inherently leads to more topologically consistent segmentation maps.
Additionally, Fig. 10 shows examples of the segmentation maps obtained by the model proposed by Galdran et al., [2022] before and after applying the proposed RR module as a post-processing step.
GT Galdran et al., [2022] Galdran et al., [2022] with RR
Refer to caption Refer to caption Refer to caption
Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to captionRefer to caption
Figure 10: Examples of segmentation maps obtained by the model of Galdran et al., [2022] before and after applying the proposed RR module as a post-processing step. A few notable differences between the resulting segmentation maps are highlighted in colored boxes. [RITE, image 11]
As evidenced by the figure, the use of this module demonstrably enhances the quality of the segmentation results. Notably, the RR module effectively addresses the issue of false positive veins in the base segmentation map by accurately reclassifying them as arteries. Additionally, it successfully bridges the gaps between arteries and veins at crossing points (represented in white), where the original segmentation predicted a low probability for one or both classes (represented in dark purple). As mentioned above, this issue is consistently observed in the outputs of other methods, with the exception of Chen et al., [2022].